I am cleaning Orbi after 6 days of continuous data acquisition. Yuanjun’s data will be uploaded in his folder in a few hours. I will start working on Gaozhong’s and Doug’s data today. All my gel-band customers: I am working on your gel bands, they should be ready for analysis by Sat morning and will be analyzed in the order they were submitted. The queue as of today is approximately 80 hours.
MALDI TOF is finally updated with the necessary software and libraries for microorganism ID. There is currently no queue.
Orbi has been busy all week trying to catch up with the long queue. Eric’s intact proteins and Gaozhong’s digests are running now. Tomorrow morning – Yuanjun’s samples. Have a nice weekend, enjoy the Fall colors while they last!
Orbi is finally working! I am running a retention time standard, after which I will resume running your samples. It took 3 calendar weeks to get Orbi back to operating condition while the Unity Lab service agreement promises a 5-day response time. Unacceptable.
I spent whole day yesterday calibrating Orbi (with little success), and the mass drift was huge today – 20 ppm error. After some troubleshooting over the phone and some digging in Orbi’s innards, I found a loose little cable. Now I must recalibrate everything again and hope this solves the problem. I will keep you posted!
Good news! A new power supply has been installed, and the instrument appears to be working. Once the FT temperature is stable, I will calibrate the instrument and start running your samples.
MALDI is in the process of getting a new software, Biotyper which will be used for microorganism identification by the Penn State Animal Diagnostics Lab. The software installation will continue as soon as some IT issues are resolved by the Bruker colleagues.
Good news: The parts for Orbi have arrived! Now we wait for the engineer.